A basketball concession stand sells sodas for $ 1.10 each and bags of popcorn for $ 0.70$ each. One​ night,32 more sodas were sold than bags of popcorn. If the concession stand made $ 127 that​ night, how many sodas and how many bags of popcorn were​ sold?
The concession stand sold how many soda and how many bags of popcorn?
John works two jobs. As a security guard he earns $ 9.75$9.75 per hour. As a landscaper he earns  $ 13.00$13.00 per hour. One week John worked a total of 3030 hours and earned $ 308.75 .$308.75. How many hours did he work at each​ job?
The number of hours John worked as a security guard is
how many hours and as a landscaper is
how mnay hours?

Answer by addingup(3677)

soda: s
popcorn: p
s = p+32
1.10s+0.70p = 127
substitute for s;
1.10(p+32)+0.70p = 127
1.80p+35.20 = 127
1.80p = 91.80
p = 51 and:
s = p+32 = 51+32 = 83
-------------------------------
Check:
51*0.70 = 35.70
83*1.10 = 91.30
Total . . 127 Correct